# Differential privacy guarantees of Markov chain Monte Carlo algorithms

### v1

@@ Appendix A Proofs and general results for Section 3
A.4 Proof of Proposition 3.10
Let ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ be two adjacent datasets. We introduce a joint process $({X^{\mathcal{D}}_{n}},X^{{\mathcal{D}}^{\prime}}_{n})_{n=1}^{N}$ such that $({X^{\mathcal{D}}_{n}})_{n=1}^{N}$ and $(X^{{\mathcal{D}}^{\prime}}_{n})_{n=1}^{N}$ are two independent Markov chains respectively with transition kernels $P_{\mathcal{D}}$ and $P_{{\mathcal{D}}^{\prime}}$, and initial distributions $\nu_{\mathcal{D}},\nu_{{\mathcal{D}}^{\prime}}$. We now define two sets, $\mathsf{B}_{1}$ and $\mathsf{B}_{2}$, that represent respectively Assumption 3.8 and Assumption 3.6:
$\mathsf{B}_{1}$ :=$:=\left\{\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\tilde{\mathcal{D}}})-\pi_{\tilde{\mathcal{D}}}(f)\right\rvert\leqslant C_{\tilde{\delta},N},\text{ for }\tilde{\mathcal{D}}={\mathcal{D}},{\mathcal{D}}^{\prime}\right\}$ (61)
@@
Now notice that $\mathsf{B}_{1}\subset\mathsf{B}_{2}$ under our assumptions, and hence $\mathbb{P}(\mathsf{B}_{2})\geqslant\mathbb{P}(\mathsf{B}_{1})$. Indeed, an application of Assumption 3.9 gives
$\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})-\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{{\mathcal{D}}^{\prime}})\right\rvert$ $\leqslant\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})-\pi_{{\mathcal{D}}}(f)\right\rvert+\left\lvert\pi_{{\mathcal{D}}}(f)-\pi_{{\mathcal{D}}^{\prime}}(f)\right\rvert+\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{{\mathcal{D}}^{\prime}})-\pi_{{\mathcal{D}}^{\prime}}(f)\right\rvert$ (63)
@@
$\mathbb{P}(\mathsf{B}_{1})$ $=\mathbb{P}\left(\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{{\mathcal{D}}})-\pi_{{\mathcal{D}}}(f)\right\rvert\leqslant C_{\delta,N}\right)\mathbb{P}\left(\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{{\mathcal{D}}^{\prime}})-\pi_{{\mathcal{D}}^{\prime}}(f)\right\rvert\leqslant C_{\tilde{\delta},N}\right).$ (65)
Hence, by Assumption 3.8
$\mathbb{P}(\mathsf{B}_{1})\geqslant(1-\tilde{\delta})^{2}=1-(2\tilde{\delta}-\tilde{\delta}^{2}).$
Therefore, we have obtained $\mathbb{P}(\mathsf{B}_{2})\geqslant 1-(2\tilde{\delta}-\tilde{\delta}^{2}).$ The result then follows by Proposition A.5.
Appendix B Proofs and auxiliary results for Section 4
@@ B.3 Girsanov’s theorem for SDEs
Here we apply Girsanov’s theorem to obtain the Radon-Nikodym derivative of the law of two SDEs with different drifts. The precise formulation we choose here is chosen so as to be able to apply the result to the diffusion (16) featured in Section 4.2. We shall assume Novikov’s condition (74).
@@ B.4 Bounds on Stochastic Integrals
In this section we present two well known bounds for exponentials of stochastic integrals.
@@
so that applying Lemma B.2 one has
@@
Let us fix adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$. Then by the assumptions and the version of Girsanov’s theorem presented in Lemma B.1, one sees that there exists a measure $\mathbb{Q}$ under which $(X^{\mathcal{D}}_{t})_{t\in[0,T]}$ is equal in law to $(X^{{\mathcal{D}}^{\prime}}_{t})_{t\in[0,T]}$ under $\mathbb{P}$, that is
@@
Then, Lemmas 4.1 and 4.2 give that DP and Rényi-DP can be obtained by controlling respectively the tails and the moments of $\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}$.
@@
Now, applying Lemma B.3 for $u_{t}=-\sqrt{\frac{\beta}{2}}h(X^{\mathcal{D}}_{\kappa(t)},\eta_{\kappa(t)})$, which under our assumptions satisfies $\lvert u_{t}\rvert\leqslant c\sqrt{\frac{\beta}{2}},$ we find
@@
Finally, Lemma 4.1 gives the wanted $(\varepsilon,\delta)$-DP bound.
Now we derive the bound on the Rényi-DP for $\alpha>1$. Lemma 4.2 shows that it is sufficient to bound
@@
We can bound the first term using that $\lvert h(X^{\mathcal{D}}_{\kappa(t)},\eta_{\kappa(t)})\rvert\leqslant c$, while the second term can be dealt applying Lemma B.2 for $u_{t}=(1-\alpha)(\beta/2)^{1/2}h(X^{\mathcal{D}}_{\kappa(t)},\eta_{\kappa(t)})$. This gives
@@
By Lemma 4.2 this implies Rényi-DP with $\varepsilon=\frac{\beta}{4}Tc^{2}\alpha.$
B.6 Proof of Proposition 4.6
@@
Now observe that by Lemma B.1, under the measure $\mathbb{Q}$ given as
@@
Now, we wish to apply Lemmas 4.1 and 4.2 to obtain the wanted guarantees respectively on the DP and Rényi-DP of our randomised algorithm. For any $\delta>0$ and $\alpha>1$, let us set $\varepsilon_{\delta}=C_{2}/4+\sqrt{C_{2}\log(1/\delta)}$ and $\varepsilon_{\alpha}=\alpha C_{2}/4$, where $C_{2}=\beta(C(L+1)+c)^{2}$. Proceeding as in the proof of Proposition 4.5 one may show that
$\mathbb{P}\biggr{(}\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\geqslant\varepsilon_{\delta}\biggr{)}\leqslant\delta,\qquad\mathsf{D}_{\alpha}(\mathbb{P}|\mathbb{Q})\leqslant\varepsilon_{\alpha},$ (110)
which is sufficient to apply Lemmas 4.1 and 4.2.
Appendix C Proofs for Section 5
C.1 Proof of Theorem 5.2
@@ Appendix C Proofs for Section 5
$\lvert e_{n}-\gamma\nabla K(x^{\mathcal{D}}_{n})-\gamma\nabla K(x^{{\mathcal{D}}^{\prime}}_{n})\rvert^{2}\leqslant(1-2\gamma\mu+\gamma^{2}L^{2})\lvert e_{n}\rvert^{2}.$ (112)
@@
$0\leqslant 1-2\gamma\mu+\gamma^{2}L^{2}\leqslant 1$ (113)
For the lower bound one can show by standard calculus that the minimum value attained for $\gamma>0$ in the above expression is $1-\mu^{2}/L^{2}$. To see that this is greater than 00, note that by the monotonicity assumption (26), applying Cauchy-Schwarz one must have
$\lvert\nabla K(x)-\nabla K(y)\rvert\geqslant\mu\lvert x-y\rvert,$ (114)
and therefore one must have $L\geqslant\mu$. The upper bound in (113) follows simply by the assumption $\gamma\in(0,2\mu/L^{2})$. Then since $\sqrt{1-x}\leqslant 1-x/2$ for $x\in[0,1]$, one may conclude
$\lvert e_{n}-\gamma\nabla K(x^{\mathcal{D}}_{n})-\gamma\nabla K(x^{{\mathcal{D}}^{\prime}}_{n})\rvert\leqslant(1-\gamma\mu+\gamma^{2}L^{2}/2)\lvert e_{n}\rvert.$ (115)
By the assumption that that $\lvert\nabla V\rvert,\lvert\nabla V^{\prime}\rvert\leqslant c$, one therefore has thatcolor=red!20]AD: for the future, see more assumption on $\nabla V^{{\mathcal{D}}}$, in fact we can deal with less restricted assumption using reflexion coupling! !-)
$\lvert e_{n+1}\rvert$ $\leqslant\lvert e_{n}-\gamma\nabla K(x^{\mathcal{D}}_{n})-\gamma\nabla K(x^{{\mathcal{D}}^{\prime}}_{n})\rvert+\lvert\gamma\nabla V^{{\mathcal{D}}}(x^{\mathcal{D}}_{n})-\gamma\nabla V^{{\mathcal{D}}^{\prime}}(x^{{\mathcal{D}}^{\prime}}_{n})\rvert$
$\leqslant(1-\gamma\mu+\gamma^{2}L^{2}/2)\lvert e_{n}\rvert+2\gamma c.$ (116)
@@
$\lvert e_{n}\rvert\leqslant\frac{2c}{\mu-\gamma L^{2}/2}.$ (117)
@@
$\sup_{t\geqslant 0}\;\lvert X^{\mathcal{D}}_{\kappa_{\gamma}(t)}-X^{{\mathcal{D}}^{\prime}}_{\kappa_{\gamma}(t)}\rvert\leqslant\frac{2c}{\mu-\gamma L^{2}/2}.$ (119)
@@
$X^{{\mathcal{D}}}_{t}-X^{{\mathcal{D}}^{\prime}}_{t}=X^{\mathcal{D}}_{\kappa_{\gamma}(t)}-X^{{\mathcal{D}}^{\prime}}_{\kappa_{\gamma}(t)}-(t-\kappa_{\gamma}(t))[\nabla U_{\mathcal{D}}(X^{\mathcal{D}}_{\kappa_{\gamma}(t)})-\nabla U_{{\mathcal{D}}^{\prime}}(X^{{\mathcal{D}}^{\prime}}_{\kappa_{\gamma}(t)})],$ (120)
@@
$\sup_{t\geqslant 0}\;\lvert X^{{\mathcal{D}}}_{t}-X^{{\mathcal{D}}^{\prime}}_{t}\rvert\leqslant\frac{2c}{\mu-\gamma L^{2}/2}.$ (122)
Therefore one sees that the assumptions of Proposition 4.6 are satisfied for $L,c>0$ as in the Proposition, and $C=\frac{2c}{\mu-\gamma L^{2}/2}$, and the result therefore follows by choosing $T=n\gamma$
C.2 Proof of Theorem 5.6
@@
$\lvert e_{n}-\frac{\gamma}{s}\sum_{i\in A_{n+1}}(\nabla k(x^{\mathcal{D}}_{n})-$ $\nabla k(x^{{\mathcal{D}}^{\prime}}_{n}))\rvert^{2}\leqslant(1-2\mu+L^{2}\gamma^{2})\lvert e_{n}\rvert^{2},$ (125)
so that since as before one has $0\leqslant 1-2\mu+L^{2}\gamma^{2}\leqslant 1$, one may square root to see that
@@
$\sup_{t\geqslant 0}\lvert Y^{\mathcal{D}}_{t}-Y^{{\mathcal{D}}^{\prime}}_{t}\rvert\leqslant\frac{2c}{\mu-\gamma L^{2}/2},$ (128)
at which point as before the result follows from Proposition 4.6, paying attention to the value of $\beta>0$ assumed in the definition (31) of the SGD algorithm.
C.3 Proof of Theorem 5.7
@@
$\lvert e_{n+1}\rvert\leqslant(1-\gamma\mu+\gamma^{2}L^{2}/2)\lvert e_{n}\rvert+\frac{2c\gamma}{s},$ (130)
@@
$\sup_{n\geqslant 1}\lvert e_{n}\rvert\leqslant\frac{2c}{s(\gamma L^{2}/2-\mu)}.$ (131)
Then extending this bound to the continuous interpolation (127) as in the proofs of Theorems 5.2 and 5.6, and applying Proposition 4.6 as before, the result follows.
@@
Therefore one may apply Proposition 4.5 to (127), and then since the law of $(Y^{\mathcal{D}}_{0},Y^{\mathcal{D}}_{\gamma},...Y^{\mathcal{D}}_{n\gamma})$ is equal to the law of $(x^{\mathcal{D}}_{0},...,x^{\mathcal{D}}_{n})$ and $(Y^{\mathcal{D}}_{0},Y^{\mathcal{D}}_{\gamma},...Y^{\mathcal{D}}_{n\gamma})$ is equal to a mapping from $(Y_{t})_{t\in[0,n\gamma]}$, the result follows from the data processing inequality (see Theorem 9 in van Erven & Harremos (2014)).

### v2

@@ Appendix A Proofs and general results for Section˜3
A.5 Proof of Proposition˜3.10
Let ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ be two adjacent datasets. We introduce a joint process $({X^{\mathcal{D}}_{n}},X^{{\mathcal{D}}^{\prime}}_{n})_{n=1}^{N}$ such that $({X^{\mathcal{D}}_{n}})_{n=1}^{N}$ and $(X^{{\mathcal{D}}^{\prime}}_{n})_{n=1}^{N}$ are two independent Markov chains respectively with transition kernels $P_{\mathcal{D}}$ and $P_{{\mathcal{D}}^{\prime}}$, and initial distributions $\nu_{\mathcal{D}},\nu_{{\mathcal{D}}^{\prime}}$. We now define two sets, $\mathsf{B}_{1}$ and $\mathsf{B}_{2}$, that represent respectively Assumption˜3.8 and Assumption˜3.6:
$\mathsf{B}_{1}$ $:=\left\{\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\tilde{\mathcal{D}}})-\pi_{\tilde{\mathcal{D}}}(f)\right\rvert\leqslant C_{\tilde{\delta},N},\text{ for }\tilde{\mathcal{D}}={\mathcal{D}},{\mathcal{D}}^{\prime}\right\}$ (67)
@@
Now notice that $\mathsf{B}_{1}\subset\mathsf{B}_{2}$ under our assumptions, and hence $\mathbb{P}(\mathsf{B}_{2})\geqslant\mathbb{P}(\mathsf{B}_{1})$. Indeed, an application of Assumption˜3.9 gives
[…]
@@
$\mathbb{P}(\mathsf{B}_{1})$ $=\mathbb{P}\left(\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{{\mathcal{D}}})-\pi_{{\mathcal{D}}}(f)\right\rvert\leqslant C_{\tilde{\delta},N}\right)\mathbb{P}\left(\left\lvert\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{{\mathcal{D}}^{\prime}})-\pi_{{\mathcal{D}}^{\prime}}(f)\right\rvert\leqslant C_{\tilde{\delta},N}\right).$ (71)
Hence, by Assumption˜3.8
$\mathbb{P}(\mathsf{B}_{1})\geqslant(1-\tilde{\delta})^{2}=1-(2\tilde{\delta}-\tilde{\delta}^{2}).$ (72)
Therefore, we have obtained $\mathbb{P}(\mathsf{B}_{2})\geqslant 1-(2\tilde{\delta}-\tilde{\delta}^{2}).$ The result then follows by Proposition˜A.6.
Appendix B Proofs and auxiliary results for Section˜4
@@ B.3 Girsanov’s theorem for SDEs
Here we apply Girsanov’s theorem to obtain the Radon-Nikodym derivative of the law of two SDEs with different drifts. The precise formulation we choose here is chosen so as to be able to apply the result to the diffusion (18) featured in Section˜4.2. We shall assume Novikov’s condition (81).
@@ B.4 Bounds on Stochastic Integrals
In this section we present two well-known bounds for exponentials of stochastic integrals.
@@
so that applying Lemma˜B.2 one has
@@
Let us fix adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$. Then by the assumptions and the version of Girsanov’s theorem presented in Lemma˜B.1, one sees that there exists a measure $\mathbb{Q}$ under which $(X^{\mathcal{D}}_{t})_{t\in[0,T]}$ is equal in law to $(X^{{\mathcal{D}}^{\prime}}_{t})_{t\in[0,T]}$ under $\mathbb{P}$, that is
@@
Then, Lemmas˜4.1 and 4.2 give that DP and Rényi-DP can be obtained by controlling respectively the tails and the moments of $\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}$.
@@
Now, applying Lemma˜B.3 for $u_{t}=-\sqrt{\frac{\beta}{2}}h(X^{\mathcal{D}}_{\kappa(t)},\eta_{\kappa(t)})$, which under our assumptions satisfies $\lvert u_{t}\rvert\leqslant c\sqrt{\frac{\beta}{2}},$ we find
@@
Finally, Lemma˜4.1 gives the wanted $(\varepsilon,\delta)$-DP bound.
Now we derive the bound on the Rényi-DP for $\alpha>1$. Lemma˜4.2 shows that it is sufficient to bound
@@
We can bound the first term using that $\lvert h(X^{\mathcal{D}}_{\kappa(t)},\eta_{\kappa(t)})\rvert\leqslant c$, while the second term can be dealt applying Lemma˜B.2 for $u_{t}=(1-\alpha)(\beta/2)^{1/2}h(X^{\mathcal{D}}_{\kappa(t)},\eta_{\kappa(t)})$. This gives
@@
By Lemma˜4.2 this implies Rényi-DP with $\varepsilon=\frac{\beta}{4}Tc^{2}\alpha.$
B.6 Proof of Proposition˜4.5
@@
Now observe that by Lemma˜B.1, under the measure $\mathbb{Q}$ given as
@@
Now, we wish to apply Lemmas˜4.1 and 4.2 to obtain the wanted guarantees respectively on the DP and Rényi-DP of our randomised algorithm. For any $\delta>0$ and $\alpha>1$, let us set $\varepsilon_{\delta}=C_{2}/4+\sqrt{C_{2}\log(1/\delta)}$ and $\varepsilon_{\alpha}=\alpha C_{2}/4$, where $C_{2}=\beta(C(L+1)+c)^{2}$. Proceeding as in the proof of Proposition˜4.6 one may show that
$\mathbb{P}\biggr(\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\geqslant\varepsilon_{\delta}\biggr)\leqslant\delta,\qquad\mathsf{D}_{\alpha}(\mathbb{P}\|\mathbb{Q})\leqslant\varepsilon_{\alpha},$ (117)
which is sufficient to apply Lemmas˜4.1 and 4.2.
Appendix C Proofs for Section˜5
C.1 Proof of Theorem˜5.2
@@ Appendix C Proofs for Section˜5
$\lvert e_{n}-\gamma\nabla K(x^{\mathcal{D}}_{n})-\gamma\nabla K(x^{{\mathcal{D}}^{\prime}}_{n})\rvert^{2}\leqslant(1-2\gamma a+\gamma^{2}L^{2})\lvert e_{n}\rvert^{2}.$ (119)
@@
$0\leqslant 1-2\gamma a+\gamma^{2}L^{2}\leqslant 1.$ (120)
For the lower bound one can show by standard calculus that the minimum value attained for $\gamma>0$ in the above expression is $1-a^{2}/L^{2}$. To see that this is greater than 0, note that by the monotonicity assumption (28), applying Cauchy-Schwarz one must have
$\lvert\nabla K(x)-\nabla K(y)\rvert\geqslant a\lvert x-y\rvert$, (121)
and therefore one must have $L\geqslant a$. The upper bound in (120) follows simply by the assumption $\gamma\in(0,2a/L^{2})$. Then since $\sqrt{1-x}\leqslant 1-x/2$ for $x\in[0,1]$, one may conclude
$\lvert e_{n}-\gamma\nabla K(x^{\mathcal{D}}_{n})-\gamma\nabla K(x^{{\mathcal{D}}^{\prime}}_{n})\rvert\leqslant(1-\gamma a+\gamma^{2}L^{2}/2)\lvert e_{n}\rvert$. (122)
By the assumption that that $\lvert\nabla V\rvert,\lvert\nabla V^{\prime}\rvert\leqslant c$, one therefore has that
[…]
$\leqslant(1-\gamma a+\gamma^{2}L^{2}/2)\lvert e_{n}\rvert+2\gamma c.$ (123)
@@
$\lvert e_{n}\rvert\leqslant\frac{2c}{a-\gamma L^{2}/2}.$ (124)
@@
$\sup_{t\geqslant 0}\;\lvert X^{\mathcal{D}}_{\kappa_{\gamma}(t)}-X^{{\mathcal{D}}^{\prime}}_{\kappa_{\gamma}(t)}\rvert\leqslant\frac{2c}{a-\gamma L^{2}/2}.$ (126)
@@
[…]
@@
$\sup_{t\geqslant 0}\;\lvert X^{{\mathcal{D}}}_{t}-X^{{\mathcal{D}}^{\prime}}_{t}\rvert\leqslant\frac{2c}{a-\gamma L^{2}/2}.$ (129)
Therefore one sees that the assumptions of Proposition 4.5 are satisfied for $L,c>0$ as in the Proposition, and $C=\frac{2c}{a-\gamma L^{2}/2}$, and the result therefore follows by choosing $T=n\gamma$
C.2 Proof of Theorem˜5.6
@@
$\lvert e_{n}-\frac{\gamma}{s}\sum_{i\in A_{n+1}}(\nabla k(x^{\mathcal{D}}_{n})$- $\nabla k(x^{{\mathcal{D}}^{\prime}}_{n}))\rvert^{2}\leqslant(1-2a+L^{2}\gamma^{2})\lvert e_{n}\rvert^{2},$ (132)
so that since as before one has $0\leqslant 1-2a+L^{2}\gamma^{2}\leqslant 1$, one may square root to see that
@@
$\sup_{t\geqslant 0}\lvert Y^{\mathcal{D}}_{t}-Y^{{\mathcal{D}}^{\prime}}_{t}\rvert\leqslant\frac{2c}{a-\gamma L^{2}/2},$ (135)
at which point as before the result follows from Proposition˜4.5, paying attention to the value of $\beta>0$ assumed in the definition (33) of the SGD algorithm.
C.3 Proof of Theorem˜5.7
@@
$\lvert e_{n+1}\rvert\leqslant(1-\gamma a+\gamma^{2}L^{2}/2)\lvert e_{n}\rvert+\frac{2c\gamma}{s},$ (137)
@@
$\sup_{n\geqslant 1}\lvert e_{n}\rvert\leqslant\frac{2c}{s(\gamma L^{2}/2-\mathcal{Y})}.$ (138)
Then extending this bound to the continuous interpolation (134) as in the proofs of Theorems˜5.2 and 5.6, and applying Proposition˜4.5 as before, the result follows.
@@
Therefore one may apply Proposition˜4.6 to (134), and then since the law of $(Y^{\mathcal{D}}_{0},Y^{\mathcal{D}}_{\gamma},...Y^{\mathcal{D}}_{n\gamma})$ is equal to the law of $(x^{\mathcal{D}}_{0},...,x^{\mathcal{D}}_{n})$ and $(Y^{\mathcal{D}}_{0},Y^{\mathcal{D}}_{\gamma},...Y^{\mathcal{D}}_{n\gamma})$ is equal to a mapping from $(Y_{t})_{t\in[0,n\gamma]}$, the result follows from the data processing inequality (see Theorem 9 in van Erven & Harremos (2014)).